# Compact: Common-token Optimized Model Pruning Across Channels and Tokens

### v1

@@ Abstract
Making LLMs more efficient in memory, latency, and serving cost is crucial for edge deployment, interactive applications, and sustainable inference at scale. Pruning is a key technique toward this goal. However, prior pruning methods are limited: width pruning often breaks the standard transformer layout or requires custom inference code, while depth pruning removes entire layers and can cause abrupt accuracy drops. In this work, we propose Compact, which jointly (i) prunes rare vocabulary to shrink embedding/unembedding and (ii) prunes FFN intermediate channels using common-token–weighted activations, aligning importance with the post-pruning token distribution. Compact enjoys merits of both depth and width pruning, such as: deployment-friendliness (keeps a standard transformer architecture), scale-adaptivity (trade off vocab vs. FFN pruning), training-free operation with competitive pruning time, and strong memory savings alongside throughput gains. Experiments across Qwen, LLaMA, and Gemma families (0.5B–70B) show state-of-the-art downstream task performance at similar or higher pruning ratios, with substantial reductions in parameters, GPU memory, and end-to-end latency.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance across a wide range of natural language tasks, but their ever-growing parameter counts—reaching billions to hundreds of billions—make deployment expensive in terms of memory, inference time, and energy cost. To broaden access and enable real-world applications such as on-device inference, classroom use, or latency-sensitive systems, it is crucial to compress LLMs while retaining as much performance as possible.
@@
Pruning has been a major line of compression work, mainly categorized into depth pruning and width pruning. Depth pruning often removes entire transformer blocks (Kim et al., 2024; Song et al., 2024; Gromov et al., 2025), while width pruning trims hidden dimensions such as FFN channels or attention heads (Ma et al., 2023; Ashkboos et al., 2024; An et al., 2024). However, these approaches are limited in two ways: (i) They prune largely mechanistically, without analyzing where parameters are concentrated within LLMs (embeddings, FFNs, or attention). This blind pruning means that methods that work for large LLMs often fail for small language models (SLMs), where parameter distributions differ. (ii) They ignore the linguistic nature of NLP models: not all tokens are equally important, yet pruning typically treats all tokens as if they contribute equally. These oversights lead to non-robust pruning performance, especially across scales.
We begin by systematically analyzing parameter distributions across model families and scales. This reveals a clear pattern: embeddings (vocabulary and unembedding layers) dominate in smaller models, while FFNs dominate in larger models. This explains why prior pruning methods lack robustness across scales—they prune the same way regardless of where redundancy actually lies. A second insight comes from the statistics of natural language: token frequencies follow a Zipfian distribution, meaning that rare tokens occur extremely infrequently and contribute little to downstream performance. Removing such rare tokens from the vocabulary reduces embedding size without significantly affecting accuracy, because language tasks are overwhelmingly driven by common tokens.
@@
These two insights naturally lead to our new approach Compact, a simple but powerful pruning framework with two modules: (i) Vocabulary pruning removes rare tokens and shrinks embedding/unembedding matrices, directly reducing parameters and memory usage, especially in SLMs. (ii) Common-token–weighted FFN pruning further reduces redundancy by scoring channels using activations, but weighting only the common tokens that remain valid after vocabulary pruning. Together, these two complementary modules address the limitations of prior work: pruning is now guided by parameter distribution, respects the linguistic structure of language tasks, and remains compatible with the standard transformer architecture.
@@
We evaluate Compact on diverse LLM families (Qwen 2.5, LLaMA 3/3.1/3.2, and Gemma 3) and across scales from 0.5B to 70B parameters. We test on seven downstream benchmarks (MMLU (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2020), ARC-C/E (Clark et al., 2018; 2018), PIQA (Bisk et al., 2020), GSM8K (Cobbe et al., 2021)) and also measure pruning time, inference throughput, and GPU memory usage. Our experiments highlight three phenomena: (i) Scale robustness: Compact maintains state-of-the-art performance at high pruning ratios even for small models, avoiding the collapse seen in prior work. (ii) Smooth degradation: Unlike depth pruning, which shows abrupt performance drops, Compact degrades gracefully with higher pruning. (iii) End-to-end efficiency: Compact’s pruning yields substantial GPU memory savings and improved throughput.
@@ 2 Related Work
Depth Pruning.
One well-explored branch of structured pruning is depth pruning (Kim et al., 2024; Song et al., 2024; He et al., 2024; Lu et al., 2024; Gromov et al., 2025; Chen et al., 2025), which removes entire transformer blocks from the model. Depth pruning is desirable because it only reduces the number of layers in an LLM while still maintaining a standard transformer architecture, allowing the pruned model to be compatible with a wide variety of existing inference frameworks. Shortened LLaMA (Kim et al., 2024) prunes the layers that minimize the increase in perplexity. SLEB (Song et al., 2024) uses a similar metric, but recalculates each layer’s importance after iteratively pruning one layer at a time. Gromov et al. (2025) prunes layers using an angular similarity metric. LLM-Streamline (Chen et al., 2025) has found that training a lightweight network to replace pruned parameters can mitigate performance loss, significantly outperforming existing methods. However, LLM-Streamline requires training the model to achieve performance improvements, which can take hours or days depending on the model size, as well as significant GPU resources. Hence, in this paper we will focus on training-free methods, which can be completed in minutes on a single GPU instead.
While depth pruning excels in its practicality, the coarse-grained removal of entire layers can lead to significant performance drops. In contrast, Compact removes only specific rows and columns of a weight matrix, offering a more fine-grained approach that improves performance.
Width Pruning.
An alternative approach to structured pruning is width pruning (Ma et al., 2023; Ashkboos et al., 2024; An et al., 2024; Xia et al., 2024; Dery et al., 2024; Gao et al., 2024b; Guo et al., 2025), which prunes each transformer layer in the LLM. There are a variety of methods to prune a transformer layer. SliceGPT (Ashkboos et al., 2024) prunes the hidden size by performing orthogonal matrix transformations and low-rank decomposition. LLM-Pruner (Ma et al., 2023) and LoRAPrune (Zhang et al., 2024) identifies and removes coupled structures using gradient information. FLAP (An et al., 2024) prunes FFN channels and self-attention heads based on a channel stability metric. Bonsai (Dery et al., 2024) builds a linear perturbation model of module importances, at the cost of significantly increased pruning time.
Recent works (Yin et al., 2024; Wang et al., 2025) have shown that non-uniform sparsity can improve pruning performance. However, these width pruning methods introduce structural irregularities that deviate from the transformer architecture. This makes it impractical to deploy width-pruned models, as they require specialized code for inference and are not compatible with standard inference libraries. Compact fixes these issues by ensuring that pruned models maintain a standard transformer architecture, enabling a deployment-friendly width-pruning method that outperforms depth pruning techniques.
@@ 3 Proposed Method: Compact
Before designing an effective pruning strategy, we first analyze where parameters are concentrated within modern decoder-only transformers. These models consist of three major groups of parameters: (i) vocab parameters, located in the embedding and unembedding layers, which map between the vocabulary space of size $V$ and the hidden dimension $D$; (ii) attention parameters, from the self-attention blocks; and (iii) FFN parameters, from the feed-forward blocks within each layer.
Formally, the embedding and unembedding layers each contain $VD$ parameters, giving
@@
for $L$ layers. For attention, the number of parameters depends on whether grouped query attention (GQA) is used (Ainslie et al., 2023). With $H$ denoting the ratio of attention heads to KV heads, the count is
@@
Figure 1: Parameter Distribution across Qwen 2.5 Models of Different Scales
Asymptotically, $N_{\text{FFN}}$ and $N_{\text{attention}}$ scale as $O(LD^{2})$, while $N_{\text{vocab}}$ scales only as $O(D)$, as $V$ is kept constant when scaling. Thus, as model size grows, vocab parameters become proportionally smaller. Conversely, for smaller models, vocab parameters can constitute a significant fraction of the total, making large vocabularies suboptimal (Tao et al., 2024). We can observe this effect empirically by calculating the relative proportions of each parameter group on popular model families. Figure 1 shows our empirical analysis on the Qwen 2.5 model family (Yang et al., 2024a), which validates our theoretical analysis. Proportions of other model families can be found in Appendix A.1. This motivates our strategy: vocabulary pruning is an efficient way to reduce parameters, especially in small-to-medium LLMs, while FFN pruning is critical for large models.
@@ 3.2 From Rare to Common: Rationale of Vocabulary Pruning
We define the set $S$ as the $V-V^{\prime}$ rarest tokens in the vocabulary. While these tokens can be directly removed—by pruning the corresponding rows in the embedding and unembedding matrices and deleting their merge rules—the key insight is conceptual: these tokens will never be generated in the pruned model. Consequently, pruning shifts the effective distribution of inputs toward common tokens. This means that subsequent optimization steps should focus on preserving performance under the common-token distribution rather than the full distribution.
This step is highly efficient: it requires no calibration data or forward passes and introduces no additional hyperparameters.
@@ 3.3 Intermediate Pruning under the Common-Token Distribution
Pruning vocabulary parameters alone reduces the embedding size but does not address redundancies in the FFNs, which dominate parameter count in large models. To prune FFNs, we adopt an activation-based criterion.
The standard act2 method (Muralidharan et al., 2024; Sandri et al., 2025) defines the importance of channel $k$ as
$I_{k}=\sum_{i=1}^{N}\bigl{(}\sigma(X_{i}W_{1})\odot X_{i}W_{2}\bigr{)}_{k}^{2},$ (4)
summing squared activations over a calibration dataset. However, this equally weights all tokens, including those in $S$.
Since tokens in $S$ will never appear in the input after pruning, their activations should not guide channel importance. We therefore introduce common act2, a weighted variant:
$I_{k}=\sum_{i=1}^{N}w_{i}\bigl{(}\sigma(X_{i}W_{1})\odot X_{i}W_{2}\bigr{)}_{k}^{2},\quad w_{i}=\begin{cases}0&x_{i}\in S,\\ 1&\text{otherwise}.\end{cases}$ (5)
This ensures that FFN pruning is explicitly optimized for the distribution of common tokens that remain valid in the pruned model.
Table 1: Advantages of Compact.
Depth pruning Width pruning Compact
ShortGPT LaCo SliceGPT 2SSP (ours)
Maintains Arch. $\checkmark$ $\checkmark$ $\checkmark$
Scale-adaptive $\checkmark$ $\checkmark$
Inference speedups $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Arch. agnostic $\checkmark$
Linguistically grounded $\checkmark$
@@ 3.4 Compact: Joint Pruning Pipeline
1:Model $M$, calibration dataset $D$, target vocabulary size $V^{\prime}$, target intermediate size $I^{\prime}$
@@
3:Run forward passes of $M$ on $D$, collect squared activations.
4:For each channel $k$, compute importance $I_{k}$ using common act2 (Eq. 5).
@@
6:  Prune $I-I^{\prime}$ least important channels (remove rows of $W_{1},W_{2}$ and columns of $W_{\text{down}}$).
@@
8:Prune vocab parameters: remove final $V-V^{\prime}$ rows of embedding and unembedding matrices; delete merges for tokens in $S$.
@@
Advantages of Compact.
i) Compact is scale-adaptive. By exposing two orthogonal knobs: (i) vocabulary pruning at the embedding/unembedding layers and (ii) common-token–weighted pruning of FFN intermediate channels, Compact can be tuned for small LLMs (emphasize vocab pruning) or large LLMs (emphasize intermediate pruning), or any mix to meet a target budget. This tunability preserves capacity on frequent tokens while enabling strong compression across model scales, all without altering the standard decoder-only architecture. ii) Compact is compatible with LLM frameworks. One weakness of most width pruning methods is that they do not maintain a standard transformer architecture. This is indeed the case with SliceGPT, which prunes the hidden size in all layers except for the final one. As a result, models pruned with SliceGPT are not compatible with the Huggingface transformers library, vLLM, or any other inference engines, limiting its practicality. In this aspect, Compact is more akin to depth pruning methods, since pruning only the vocabulary and intermediate size does not affect the transformer architecture. As a result, Compact models are compatible with all inference engines, making it a practical approach to width pruning. The full advantages of Compact is summarized in Table 1.

### v2

@@ Abstract
Making large language models (LLMs) more efficient in memory, latency, and serving cost is crucial for edge deployment, interactive applications, and sustainable inference at scale. Pruning is a promising technique, but existing pruning methods are limited: width pruning often breaks the standard transformer layout, requiring custom inference code, while depth pruning can cause abrupt accuracy drops. Also, while many pruning approaches are effective against LLMs, they struggle to maintain performance on small language models (SLMs). In this work, we propose Compact, which jointly (i) prunes rare vocabulary to shrink embedding/LM head layers and (ii) prunes FFN intermediate channels using common-token–weighted activations, aligning importance with the post-pruning token distribution. Compact inherits strengths of both depth and width pruning, such as: deployment-friendliness (keeps a standard transformer architecture), scale-adaptivity (trade off vocab. vs. FFN pruning), competitive pruning times, and strong memory savings alongside throughput gains. Experiments across Qwen, LLaMA, and Gemma families (0.5B–70B) show state-of-the-art downstream performance, with substantial reductions in parameters, GPU memory, and latency [^1] All code will be released, and the method will be packaged as a plug-in tool..
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance across a wide range of natural language tasks, but their ever-growing parameter counts, reaching billions to hundreds of billions, make deployment expensive in terms of memory, inference time, and energy cost. To broaden access and enable real-world applications such as on-device inference, classroom use, or latency-sensitive systems, it is crucial to compress LLMs while retaining as much performance as possible.
@@
Quantization (Frantar et al., 2022; Lin et al., 2024) and pruning (Frantar & Alistarh, 2023; Sun et al., 2024) have been a major line of compression work. This work focuses on structured pruning, removing entire rows and columns of weight matrices. Structured pruning is mainly categorized into depth pruning and width pruning. Depth pruning removes entire transformer blocks (Kim et al., 2024; Song et al., 2024; Gromov et al., 2025), but the coarse-grained removal of layers leads to sharp performance drops. Width pruning trims hidden dimensions such as FFN channels or attention heads (Ma et al., 2023; Ashkboos et al., 2024; An et al., 2024), but they typically deviate from a standard transformer architecture and require custom inference code. In addition, these approaches are limited in three other ways: (i) They prune largely mechanistically, without analyzing where parameters are concentrated within LLMs (embeddings, FFNs, or attention). This blind pruning means that methods that work for large LLMs often fail for SLMs, as they have different parameter distributions. (ii) They ignore the linguistic nature of NLP models: not all tokens are equally important, yet pruning typically treats all tokens as if they contribute equally. (iii) They often require custom implementation changes to accommodate every model family, making implementation maintenance tedious. These oversights lead to non-robust pruning performance across scales.
@@
To address these issues, we propose Compact, a simple but powerful pruning framework with two modules: (i) Vocabulary pruning removes rare tokens and shrinks embedding/LM head matrices, directly reducing parameters and memory usage, especially in SLMs. (ii) Common-token–weighted FFN pruning further reduces redundancy by scoring channels using activations, but weighting only the common tokens that remain valid after vocabulary pruning. Together, these two complementary modules address the limitations of prior work: pruning is now guided by parameter distribution, respects the linguistic structure of language tasks, remains compatible with existing inference frameworks, and is architecture-agnostic across most model families.
We systematically analyze parameter distributions across model families and scales. This reveals a clear pattern: embeddings (vocabulary and LM head layers) are important in SLMs, while FFNs dominate in larger models. This explains why prior pruning methods lack robustness across scales—they prune the same way regardless of where redundancy actually lies. A second insight comes from the statistics of natural language: token frequencies follow a Zipfian distribution (Zhemchuzhina et al., 2022), meaning that rare tokens occur extremely infrequently and contribute little to downstream performance. Removing such rare tokens from the vocabulary reduces embedding size without significantly affecting performance, because language tasks are overwhelmingly driven by common tokens. Together, these observations validate the effectiveness of the Compact method.
@@
We evaluate Compact on diverse LLM families (Qwen 2.5, LLaMA 3.1/3.2, and Gemma 3) and across scales from 0.5B to 70B parameters. We test on seven downstream benchmarks (MMLU (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2020), ARC-C/E (Clark et al., 2018), PIQA (Bisk et al., 2020), GSM8K (Cobbe et al., 2021)) and also measure pruning time, inference throughput, and GPU memory usage. Our experiments highlight three phenomena: (i) Scale robustness: Compact maintains state-of-the-art performance at high pruning ratios even for SLMs. (ii) Smooth degradation: Unlike depth pruning, which shows abrupt performance drops, Compact degrades gracefully with higher pruning. (iii) End-to-end efficiency: Compact yields substantial GPU memory savings and improved throughput.
@@
Table 1: Advantages of Compact.
Depth pruning Width pruning Compact
ShortGPT LaCo LLM-Streamline SliceGPT 2SSP FLAP (ours)
Maintains architecture $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Scale-adaptive $\checkmark$ $\checkmark$
Inference speedups $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Fast pruning $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Architecture-agnostic $\checkmark$
Linguistically grounded $\checkmark$
@@ 2 Related Work
Depth Pruning removes entire transformer blocks while preserving the standard architecture and compatibility with common inference frameworks (He et al., 2024; Lu et al., 2024). Representative methods include Shortened LLaMA (perplexity-minimizing), SLEB (iterative recalibration), and angular-similarity pruning (Kim et al., 2024; Song et al., 2024; Gromov et al., 2025). LLM-Streamline trains a lightweight network to recover accuracy but requires hours–days and significant GPUs (Chen et al., 2025). We therefore focus on training-free pruning that runs in minutes on a single GPU; Compact can optionally be fine-tuned and outperforms training-based baselines. Because depth pruning is coarse-grained and can cause sharp drops, Compact instead prunes rows/columns for a finer-grained alternative.
Width Pruning removes hidden dimensions or channels in each layer (Xia et al., 2024; Gao et al., 2024b; Guo et al., 2025). Methods include LLM-Pruner/LoRAPrune (gradient-based) (Ma et al., 2023; Zhang et al., 2024), SliceGPT (orthogonal transforms + low-rank) (Ashkboos et al., 2024), FLAP (stability-based) (An et al., 2024), and Bonsai (perturbation modeling) (Dery et al., 2024). While effective, they often break the standard transformer layout, requiring custom inference code and limiting deployment. Compact avoids these issues by preserving architecture, yielding a deployment-friendly width-pruning method that outperforms depth pruning.
Vocabulary Size/Pruning.
The vocabulary size of a model is the number of tokens that the model can recognize. Modern LLM vocabularies often reach into the hundreds of thousands and typically remain the same across model scales within a family (Tao et al., 2024). Research into vocabulary size scaling (Tao et al., 2024) has shown that the optimal vocabulary size increases with increasing LLM size, contradicting the common practice of keeping vocabulary size constant over a wide range of model sizes. Prior work prunes vocabularies to tailor vocabulary to a target language/domain (Ushio et al., 2023b; Dorkin et al., 2025; Ushio et al., 2023a; Bogoychev et al., 2024; Yang et al., 2022; 2024b). Others prune the drafter’s LM head for speculative decoding speedups (Goel et al., 2025), which does not compress the base model used at inference. In contrast, we (i) perform non-domain-specific vocabulary pruning for general-purpose LLMs, and (ii) couple it with common-token–weighted FFN pruning, so channel scores reflect the token distribution after vocab removal. This keeps a standard Transformer layout, is training-free, and proves robust from 0.5B to 70B.
@@ 3 Proposed Method: Compact
Before designing an effective pruning strategy (Sections 3.2-3.4 ), we first analyze where parameters are concentrated within modern decoder-only transformers (Section 3.1).
@@
Figure 1: Parameter distribution across Qwen 2.5 models of different scales.
Mainstream generative LLMs consist of three major groups of parameters: (i) vocab parameters, located in the embedding and LM head layers; (ii) attention parameters, from the self-attention blocks; and (iii) FFN parameters, from the feed-forward blocks.
Formally, the embedding and LM head layers map between the vocabulary space of size $V$ and the hidden dimension $D$, giving
@@
for $L$ layers. For attention, the number of parameters depends on whether grouped query attention is used (Ainslie et al., 2023). With $H$ denoting the ratio of attention heads to KV heads, the count is
@@
Asymptotically, $N_{\text{FFN}}$ and $N_{\text{attention}}$ scale as $O(LD^{2})$$I\approx O(D)$—while $N_{\text{vocab}}$ scales only as $O(D)$, as $V$ is kept constant when scaling. Thus, as model size grows, vocab parameters become proportionally smaller. Conversely, for smaller models, vocab parameters can constitute a significant fraction of the total. We observe this empirically by calculating the relative proportions of each parameter group on popular model families. Figure 1 shows our empirical analysis on the Qwen 2.5 model family (Yang et al., 2024a), which validates our theoretical analysis. Proportions of other model families can be found in Appendix A.1. This motivates our strategy: vocabulary pruning is an efficient way to reduce parameters, especially in small-to-medium LLMs, while FFN pruning is critical for large models.
@@ 3.2 From Rare to Common: Rationale of Vocabulary Pruning
We define the set $S$ as the $V-V^{\prime}$ rarest tokens in the vocabulary. These tokens can be directly removed by pruning the corresponding rows in the embedding/LM head matrices and deleting the corresponding merge rules from the tokenizer. The key insight is conceptual: the deleted tokens will never be generated in the pruned model. This means that subsequent optimization steps should focus on preserving performance under the common-token distribution rather than the full distribution. Vocab-pruning is highly efficient: it requires no calibration data or forward passes.
@@ 3.3 Intermediate Pruning under the Common-Token Distribution
Pruning vocabulary parameters alone reduces the embedding size but does not address redundancies in the FFNs, which dominate parameter count in large models. To prune FFNs, we adopt an activation-based criterion. The standard act2 method (Muralidharan et al., 2024; Sandri et al., 2025) defines the importance of FFN intermediate channel $k$ as
$\mathcal{I}_{k}=\sum_{i=1}^{N}\bigl(\text{SiLU}(X_{i}W_{\text{gate}})X_{i}W_{\text{up}}\bigr)_{k}^{2},$ (4)
summing squared activations over a calibration dataset. Here, $X_{i}$ is FFN input and $W_{\text{gate}},W_{\text{up}}$ are model weights. However, this equally weights all tokens $x_{i}$, including $x_{i}\in S$. Since such tokens will never appear in the input after pruning, their activations should not guide channel importance. We therefore introduce common act2, a weighted variant:
$\mathcal{I}_{k}=\sum_{i=1}^{N}w_{i}\bigl(\text{SiLU}(X_{i}W_{\text{gate}})X_{i}W_{\text{up}}\bigr)_{k}^{2},\quad w_{i}=\begin{cases}0&x_{i}\in S,\\ 1&\text{otherwise}.\end{cases}$ (5)
This ensures that FFN pruning is explicitly optimized for the tokens that remain valid after pruning.
@@ 3.4 Compact: Joint Pruning Pipeline
1:Model $M$, calibration dataset $\mathcal{D}$, target vocabulary size $V^{\prime}$, target intermediate size $I^{\prime}$
@@
3:Run forward passes of $M$ on $\mathcal{D}$, collect squared activations.
4:For each channel $k$, compute importance $\mathcal{I}_{k}$ using common act2 (Eq. 5).
@@
6:  Prune $I-I^{\prime}$ least important channels (remove rows of $W_{\text{gate}},W_{\text{up}}$ and columns of $W_{\text{down}}$).
@@
8:Prune vocab parameters: remove final $V-V^{\prime}$ rows of embedding and LM head matrices; delete tokenizer merges for tokens in $S$.
@@
Advantages of Compact. i) Compact is scale-adaptive. Compact uses two different knobs for pruning: (i) vocabulary pruning at the embedding/LM head layers and (ii) common-token–weighted pruning of FFN intermediate channels. These two knobs are orthogonal, allowing Compact to be tunable for SLMs (emphasize vocab pruning) LLMs (emphasize intermediate pruning), or any mix to meet a target budget. This tunability preserves capacity on frequent tokens while enabling strong compression across model scales. ii) Compact is compatible with LLM frameworks. One weakness of most width pruning methods is that they do not maintain a standard transformer architecture. This is indeed the case with SliceGPT, which prunes the hidden size in all layers except for the final one, as well as 2SSP, which prunes entire attention modules. As a result, these methods are not compatible with the transformers library, vLLM, or any other inference engines, limiting practicality. In this aspect, Compact is similar to depth pruning, since pruning the vocabulary and intermediate size does not affect the transformer architecture. As a result, Compact models are compatible with all inference engines, making it a practical approach to width pruning. The full advantages of Compact is summarized in Table 1.